OpenStudy (anonymous):

Patty's Pet Company charges $0.55 per pound to ship pet supplies. Part A: Write an equation to determine the total cost, c, to ship p pounds of pet supplies. Use your equation to determine the cost of shipping 2 pounds of pet supplies. Part B: If the company reduces the cost to ship pet supplies by 0.05 per pound, write an equation to determine the total cost, c, to ship p pounds of pet supplies with the reduced cost

OpenStudy (anonymous):

can i get help plz

OpenStudy (anonymous):

Can I get step by step help

OpenStudy (anonymous):

@jim_thompson5910 can you help me

jimthompson5910 (jim_thompson5910):

"Patty's Pet Company charges $0.55 per pound to ship pet supplies." how much does it cost to ship exactly 1 pound of pet supplies?

OpenStudy (anonymous):

.26

jimthompson5910 (jim_thompson5910):

how are you getting that?

OpenStudy (anonymous):

.55/2

jimthompson5910 (jim_thompson5910):

"$0.55 per pound" means "$0.55 for each pound" or "$0.55 for one pound"

OpenStudy (anonymous):

.55*2 =1.10

jimthompson5910 (jim_thompson5910):

1 pound simply costs $0.55 you're overthinking

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

how about 7 pounds? how much does that cost?

OpenStudy (anonymous):

3.85

jimthompson5910 (jim_thompson5910):

and how did you get that?

OpenStudy (anonymous):

.55*7

jimthompson5910 (jim_thompson5910):

so in general, if we don't know the amount of pounds, we can just say it's p so the general cost is 0.55*p agreed?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

so that's why the cost equation is c = 0.55*p

jimthompson5910 (jim_thompson5910):

"Use your equation to determine the cost of shipping 2 pounds of pet supplies. " now plug in p = 2 into that cost equation

OpenStudy (anonymous):

c=.55*2 = 1.10

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

Part B would be c = .50*2?

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

they say nothing about p being 2 though

OpenStudy (anonymous):

c=.50*p

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

Don't i need to work it out are just leave it as the equation ?

OpenStudy (anonymous):

.50*1.10?

jimthompson5910 (jim_thompson5910):

they just want c = 0.50p for part B

OpenStudy (anonymous):

got it thanks

jimthompson5910 (jim_thompson5910):

no problem